if two varieties are tested in 5 different locations for their yeld proformance in (q/ha) with the follo results.Find out which variety which is more consistent over the location by comparing their CV value?


Variety A -12,8,10,11,14


Variety B - 15,9,7,12,12

1
Expert's answer
2022-03-14T19:32:01-0400

"CV=\frac{\sigma}{\mu}*100" %

Variety A

Mean "=\frac{12+8+10+11+14}{5}=11"

Standard deviation "=\sqrt{\frac{∑x^2}{N}-\mu^2}=\sqrt{\frac{625}{5}-11^2}=\sqrt{4}=2"

CV "=\frac{\sigma}{\mu}*100" % "=\frac{2}{11}*100" % "=18.18" %


Variety B

Mean "=\frac{15+9+7+12+12}{5}=11"

Standard deviation "=\sqrt{\frac{∑x^2}{N}-\mu^2}=\sqrt{\frac{643}{5}-11^2}=\sqrt{7.6}=2.76"

CV "=\frac{2.76}{11}*100" % "=25.09" %


Variety A is more consistent as it has the least CV value.